# Predictive Performance Comparison of Decision Policies Under Confounding

### v1

@@ 1 Introduction
Let the status quo policy consist of human-only decisions (e.g., pre-trial release decisions (Kleinberg et al., 2018), medical testing decisions (Mullainathan & Obermeyer, 2019)) made using covariates and unobserved contextual information. We would like to evaluate whether a model recommending actions by thresholding risk predictions would constitute an improvement over the current human-only policy (e.g., (Rambachan et al., 2021; Chouldechova et al., 2018)).
@@
Let the status quo policy consist of humans making decisions with the support of a predictive model, covariates, and unobserved contextual information. We would like to assess whether decisions made using the predictive model alone would yield a performance improvement over the human-algorithm combination (e.g., (Cheng et al., 2022)).
@@
Both of these examples can be formalized as confounded off-policy evaluation in the contextual bandits setting (Jung et al., 2020; Swaminathan & Joachims, 2015; Tennenholtz et al., 2021; Rambachan et al., 2022). Given observational data collected under a confounded status quo policy, our goal is to compare predictive performance against an alternative policy which would select different actions.
However, existing off-policy evaluation frameworks do not support comparison of predictive performance metrics in the settings outlined above. First, existing approaches target the expected potential outcome given actions selected under the new policy, i.e. $V(\pi)=\mathbb{E}[Y(\pi)]$ (Kallus & Zhou, 2020; Namkoong et al., 2020; Zhang et al., 2021; Uehara et al., 2022; Si et al., 2020; Kallus & Zhou, 2020; Hatt et al., 2022; Ek et al., 2023; Tennenholtz et al., 2021; Ishikawa & He, 2023). Yet real-world model evaluations require fine-grained assessments of predictive performance metrics, such as the positive and negative class performance. These measures are challenging to estimate from off-policy data because they require conditioning on a partially observed potential outcome. While Rambachan et al. (2022) develop an evaluation framework targeting predictive performance measures under confounding, their approach does not support comparisons against the status quo data-generating policy. Additionally, prior frameworks for off-policy learning and evaluation rely upon a specific causal assumption (e.g., Rosenbaum’s $\Gamma$ model (Namkoong et al., 2020; Zhang et al., 2021), instrumental variables (Lakkaraju et al., 2017; Kleinberg et al., 2018; Chen et al., 2023), or the marginal sensitivity model (MSM) (Kallus & Zhou, 2021)). In real-world auditing settings, it is critical that evaluation frameworks are compatible with a broad set of causal assumptions to support flexibility.
In this work, we develop the first framework supporting predictive-performance comparisons of decision-making policies under confounding. Our approach partially identifies the performance of a proposed policy against a status quo data-generating policy. We decompose performance measures into sets of identified and partially identified comparison statistics. We show that a diverse set of causal assumptions studied in prior off-policy evaluation literature imply uncertainty sets over partially identified sufficient statistics. We develop an approach for tightly characterizing policy regret bounds by isolating uncertainty relevant to the policy comparison. Our approach leverages the key insight that we can eliminate redundant uncertainty which does not contribute to differential policy performance (Figure 1).
@@
We propose a flexible algorithm for finite sample estimation of regret bounds under no parametric assumptions on the data-generating policy. While this procedure inherits slow non-parametric convergence rates of regression functions used to estimate bounds on comparison statistics, we also show how to construct assumption-tailored doubly robust estimators which attain fast root-$n$ convergence rates under no parametric assumptions. We conduct synthetic experiments verifying the coverage of our regret intervals. We conclude by illustrating how our framework can be used to support a pre-deployment evaluation of a proposed modification to a healthcare enrollment policy (Obermeyer et al., 2019). We make the following main contributions:
@@
  • We formulate the problem of comparative predictive performance evaluations for decision making policies under uncertainty (§$\S$ 3).
@@
  • We validate our framework theoretically and via synthetic data experiments (§$\S$ 7). We demonstrate how our framework can support a pre-deployment evaluation of a proposed healthcare enrollment policy (§$\S$ 8). We present all proofs in the Appendix.
@@ 2 Related work
Off-policy evaluation (OPE) is a widely studied problem in reinforcement learning (Uehara et al., 2022; Jiang & Li, 2016; Precup, 2000; Namkoong et al., 2020; Kallus & Zhou, 2020), contextual bandits (Si et al., 2020; Tennenholtz et al., 2021; Swaminathan & Joachims, 2015; Xu et al., 2021; Dudík et al., 2014), and econometrics (Kitagawa & Tetenov, 2018; Huber, 2019) literature. To our knowledge, we offer the first off-policy evaluation framework supporting predictive performance comparisons of decision policies.
We study the confounded offline contextual bandits setting (Jung et al., 2020; Swaminathan & Joachims, 2015; Tennenholtz et al., 2021), in which observational data is collected under a behavior policy $\pi_{0}(X_{i},U_{i})$ which acts on covariates $X_{i}$ and unmeasured confounders $U_{i}$. Given a new policy $\pi(X_{i})$ [^1] Work in the confounded offline setting typically studies updated policies that only observe covariates. When $\pi(X_{i})$ also observes $U_{i}$ regret intervals tend to be vacuous. , prior work typically estimates its value function $V(\pi)=\mathbb{E}[Y(\pi(X))]$, which is the mean potential outcome under actions selected by $\pi$. Existing off-policy evaluation methods target this value function rather than fine-grained predictive performance measures (e.g., TNR, PPV) of interest in real-world auditing settings (Kallus & Zhou, 2020; Namkoong et al., 2020; Zhang et al., 2021; Uehara et al., 2022; Si et al., 2020; Kallus & Zhou, 2020; Hatt et al., 2022; Ek et al., 2023; Tennenholtz et al., 2021; Ishikawa & He, 2023). Off-policy evaluation of predictive performance measures is challenging because it requires conditioning on a partially-observed potential outcome. Rambachan et al. (2022) provide a solution by developing a framework for robust learning and auditing of predictive models under confounding. This framework assesses the predictive performance of a probabilistic score function, such as a risk assessment. However, the score function is typically unknown for the data-generating policy $\pi_{0}$ in settings where the policy involves human decisions. In contrast, our approach directly supports predictive performance comparisons of $\pi$ against $\pi_{0}$. We show that our approach yields tighter regret intervals than would follow from extending Rambachan et al. (2022) to bound status quo policy performance $V(\pi_{0})$, then computing the regret by taking a difference $V(\pi)-V(\pi_{0})$.
@@
Our work also builds upon prior literature studying policy learning and selection under confounding (Zhang et al., 2021; Namkoong et al., 2020; Konyushova et al., 2021; Yang et al., 2022; Kuzborskij et al., 2021; Zhang et al., 2021; Kallus & Zhou, 2021, 2018; Hatt et al., 2022; Ek et al., 2023; Gao & Yin, 2023; Balachandar et al., 2023). Zhang et al. (2021) develop a framework for ranking individualized treatment policies under Rosenbaum’s $\Gamma$ sensitivity model. Namkoong et al. (2020) also use Rosenbaum’s $\Gamma$ sensitivity model to partially identify $V(\pi)$ in the sequential setting. Kallus & Zhou (2021, 2018) develop an approach for learning minimax optimal decision-policies under the MSM. In principle, confounding-robust policy learning frameworks such as these can also be used to recover regret intervals against the confounded data-generating policy (Hatt et al., 2022; Ek et al., 2023; Gao & Yin, 2023). However, these techniques (1) target the mean potential outcome rather than predictive performance measures, and (2) are tied to a fixed causal assumption which may not be applicable in all contexts. In contrast, our framework is interoperable with a range of assumptions (e.g., instrumental variable, MSM, proximal variable) studied in prior OPE literature.
Kleinberg et al. (2018) develop a framework for comparing human decisions and algorithmic predictions under confounding. The contraction technique underpinning this approach compares the failure rate of judicial decisions against a predictive model under an assumption that judges have heterogeneous selection rates (Lakkaraju et al., 2017). Later developments formalize this approach into an instrumental variable framework (Chen et al., 2023; Rambachan et al., 2021). Most recently, Rambachan et al. (2022) propose a Mean Outcome Sensitivity Model (MOSM) which is implied by Rosenbaum’s $\Gamma$ sensitivity model, the MSM, and IVs. We extend this framework to our policy comparison setting by developing an approach for mapping causal assumptions to uncertainty sets around partially identified comparison statistics (see Appendix B).
@@ 3 Preliminaries
Let $\pi_{0}:\mathcal{X}\times\mathcal{U}\rightarrow\mathcal{A}$ be a status quo decision-making policy assigning actions given measured covariates $X\in\mathcal{X}$ and unmeasured confounders $U\in\mathcal{U}$. Let $\pi:\mathcal{X}\rightarrow\mathcal{A}$ be a proposed policy that assigns actions only via covariates, where $\mathcal{A}\in\{0,1\}$ is binary for both policies. [^2] Our framework is compatible with new policies which are stochastic and deterministic. However we assume that $\pi$ is known.. For example, in a medical testing context, $\pi_{0}$ is an existing (e.g., physician based) testing policy, while $\pi$ is a proposed algorithmic policy. Let $D^{\pi_{0}}$, $T^{\pi}$ be random variables indicating actions selected under the status quo and proposed policies, respectively. [^3] We sometimes omit policy superscripts over random variables to ease notation. Let $Y(1)\in\mathcal{Y}$ be the outcome of interest to the policy comparison. This is the potential outcome which would be observed given a positive decision ($D^{\pi_{0}}=1$) selected under the status quo policy (Rubin, 2005). For example, $Y(1)$ denotes the disease status of a patient, or repayment if a lendee is granted a loan. Let $Y\in\mathcal{Y}$ be the outcome in the observational data. In our selective labels setting (Lakkaraju et al., 2017), $Y$ is only observed when an instance received a positive decision ($D^{\pi_{0}}=1$) under $\pi_{0}$. We make the following standard consistency assumption on $Y$ (Rubin, 2005).
@@ 3.1 Problem formulation
3.2 Partial identification of policy performance measures
@@ 4 Regret bound identification
We now introduce our novel approach for comparing policies under confounding. Our approach yields informative regret intervals by eliminating redundant uncertainty that does not contribute to differential policy performance. We first focus on asymptotic performance bounds, and will provide finite-sample analyses in Section 6.
@@
Because $v_{0}(0,0)\in\mathbf{v}_{0}$ cancels when taking a difference across polices, this term does not contribute uncertainty to the $\delta$-regret interval, as it would with the baseline interval. We now provide a theoretical result characterizing the improvement offered by the $\delta$-interval more generally. We let $I_{\delta}(m,\mathcal{V})=\overline{R}_{\delta}(\pi,\pi_{0};m,\mathcal{V})-\underline{R}_{\delta}(\pi,\pi_{0};m,\mathcal{V})$ and $I(m,\mathcal{V})=\overline{R}(\pi,\pi_{0};m,\mathcal{V})-\underline{R}(\pi,\pi_{0};m,\mathcal{V})$ be the length of the $\delta$-regret and baseline regret intervals, respectively.
@@ 5 Mapping causal assumptions to informative regret bounds
The regret interval we recover without imposing additional causal assumptions tends to be uninformative in many settings. Therefore, we now introduce a pointwise bounding functions assumption which we use to tighten regret intervals by shrinking the uncertainty set around partially-identified $v$-statistics. We show that this assumption is implied by a range of causal assumptions studied in prior OPE literature (e.g., MSM, Rosenbaum’s $\Gamma$, IV) in Appendix B. Thus, our $\delta$-regret interval yields tighter uncertainty quantification than the baseline interval for any causal assumption implying pointwise bounding functions (Figure 2).
@@
We only use intervals $\mathcal{H}(v_{1}(t,0);\tau)$ for this set definition because an analogous bound is implied over $v_{0}(t,0)$ by the constraint requiring that $\sum_{y}v_{y}(t,0)=\rho_{t0}$. We quantify the size of $\mathcal{V}(p;\tau)$ via the Lebesgue measure, which is defined as a cartesion product over partially identified intervals $\lambda(\mathcal{V}(p;\tau))=|\overline{v}_{1}(1,0)-\underline{v}_{1}(1,0)||\overline{v}_{1}(0,0)-\underline{v}_{1}(0,0)|$. We now show that $\mathcal{V}(p;\tau)$ is essentially the smallest uncertainty set one could construct based on the observational distribution and the bounding functions.

### v2

@@ 1 Introduction
Let the status quo policy consist of human-only decisions (e.g., pre-trial release decisions (Kleinberg et al., 2018), medical testing decisions (Mullainathan & Obermeyer, 2019)) made using covariates and unobserved contextual information. We would like to evaluate whether a model recommending actions by thresholding risk predictions would constitute an improvement over the current human-only policy (e.g., Rambachan et al. (2021); Chouldechova et al. (2018)).
@@
Let the status quo policy consist of humans making decisions with the support of a predictive model, covariates, and unobserved contextual information. We would like to assess whether decisions made using the predictive model alone would yield a performance improvement over the human-algorithm combination (e.g., Cheng et al. (2022)).
@@
Both of these examples can be formalized as confounded off-policy evaluation in the contextual bandits setting (Jung et al., 2020; Swaminathan & Joachims, 2015; Tennenholtz et al., 2021; Rambachan et al., 2022). Given observational data collected under a confounded status quo policy, our goal is to compare predictive performance against an alternative policy which would assign different actions. A common strategy for off-policy evaluation under confounding involves partial identification of policy performance within an uncertainty interval (Kallus & Zhou, 2021; Rambachan et al., 2021, 2022; Pu & Zhang, 2021; Zhang et al., 2021). Yet constructing informative performance intervals is challenging because it often requires imposing untestable assumptions on the structure and magnitude of unmeasured confounders impacting the status quo policy.
In this work, we develop a novel partial identification method which reduces uncertainty in policy performance comparisons by isolating confounding-related uncertainty in the disagreement region of the action space (Figure 1). Our uncertainty cancellation approach yields tighter regret intervals than existing frameworks designed for non-comparative performance evaluations (Rambachan et al., 2022; Namkoong et al., 2020), which account for redundant uncertainty in the agreement region of the action space. Our simple approach operates by (1) decomposing policy performance measures into sets of identified and partially identified comparison statistics, then (2) cancelling common partially identified terms in the regret estimand.
We show that a diverse set of causal assumptions studied in prior off-policy evaluation literature imply uncertainty sets over partially identified comparison statistics. As a result, while existing off-policy evaluation frameworks are often tied to a fixed causal assumption (e.g., the marginal sensitivity model (Tan, 2006)) that may be unjustified in some contexts, ours is interoperable with a broad family of modern identification strategies (e.g., Rosenbaum’s $\Gamma$ model (Namkoong et al., 2020; Zhang et al., 2021), instrumental variable (Lakkaraju et al., 2017; Kleinberg et al., 2018), proximal variable (Ghassami et al., 2023), or the marginal sensitivity model (MSM) (Kallus & Zhou, 2021)).
@@
We propose a flexible plug-in algorithm for finite sample estimation of regret bounds under no parametric assumptions on the data-generating policy. While this procedure inherits slow non-parametric convergence rates of regression functions used to estimate bounds on comparison statistics, we also show how to construct assumption-tailored doubly robust estimators which attain fast root-$n$ convergence rates under no parametric assumptions. We conduct synthetic experiments verifying the coverage of our regret intervals. We conclude by illustrating how our framework can be used to support a pre-deployment evaluation of a proposed modification to a healthcare enrollment policy (Obermeyer et al., 2019). Our results demonstrate that, in some cases, our improved regret interval supports more conclusive pre-deployment assessments of decision policies than would be possible via existing off-policy evaluation approaches.
We make the following main contributions:
@@
  • We formulate the problem of comparative predictive performance evaluations for decision making policies under unmeasured confounding (§$\S$ 3).
@@
  • We validate our framework theoretically and via synthetic data experiments (§$\S$ 7). We demonstrate how our framework can support a pre-deployment evaluation of a proposed healthcare enrollment policy (§$\S$ 8). We present all proofs in the Appendix. [^1] All code for experiments is publicly available here.
@@ 2 Related Work
Off-policy evaluation (OPE) is a widely studied problem in reinforcement learning (Uehara et al., 2022; Jiang & Li, 2016; Precup, 2000; Namkoong et al., 2020; Kallus & Zhou, 2020), contextual bandits (Si et al., 2020; Tennenholtz et al., 2021; Swaminathan & Joachims, 2015; Xu et al., 2021; Dudık et al., 2014), and econometrics (Kitagawa & Tetenov, 2018; Huber, 2021) literature. We study the confounded offline contextual bandits setting (Jung et al., 2020; Swaminathan & Joachims, 2015; Tennenholtz et al., 2021), in which observational data is collected under a behavior policy $\pi_{0}(X_{i},U_{i})$ which acts on covariates $X_{i}$ and unmeasured confounders $U_{i}$.
Existing OPE frameworks typically estimate the value function $V(\pi)=\mathbb{E}[Y(\pi(X))]$ of a new policy $\pi(X_{i})$ via observational data (Kallus & Zhou, 2020; Namkoong et al., 2020; Zhang et al., 2021; Uehara et al., 2022; Si et al., 2020; Kallus & Zhou, 2020; Hatt et al., 2022; Ek et al., 2023; Tennenholtz et al., 2021; Ishikawa & He, 2023). [^2] Work in the confounded offline setting typically studies updated policies that only observe covariates. When $\pi(X_{i})$ also observes $U_{i}$ regret intervals tend to be vacuous. In this work, we instead target predictive performance measures (e.g., Accuracy, TNR, PPV), which are often of interest during policy evaluation of algorithmic risk assessments. While OPE of predictive performance measures introduces estimand-specific challenges (i.e., conditioning on a partially-observed potential outcome), approaches developed for OPE of value functions are also applicable in the algorithmic decision-making context (e.g., doubly robust estimation, reweighting). [^3] In the algorithmic decision-making context, the term selective labels is often used to describe a setting in which a potential outcome is only observed under one of the possible actions (Lakkaraju et al., 2017; Wei, 2021; Coston et al., 2021; De-Arteaga et al., 2018).
In particular, Rambachan et al. (2022) develop a partial identification framework for robust learning and evaluation of algorithmic risk assessments under confounding. This framework leverages doubly-robust approaches for OPE of predictive performance measures. One natural strategy for extending this framework to the policy comparison setting involves computing bounds on $V(\pi)$ and $V(\pi_{0})$ independently before computing the regret interval by taking a difference. However, we show that this approach yields an overly conservative regret interval by introducing redundant uncertainty in the agreement region of the policy action space. We introduce an alternative $\delta$-regret interval, which we show yields more informative bounds than this baseline approach. Moreover, we extend the Mean Outcome Sensitivity Model (MOSM) introduced by Rambachan et al. (2022) to develop uncertainty sets around partially identified policy comparison statistics which are consistent with a range of modern causal assumptions (e.g., instrumental variable, MSM, Rosenbaum’s $\Gamma$ sensitivity model).
@@
Our work also builds upon prior literature studying policy learning and selection under confounding (Zhang et al., 2021; Konyushova et al., 2021; Yang et al., 2022; Kuzborskij et al., 2021; Zhang et al., 2021; Kallus & Zhou, 2021, 2018; Hatt et al., 2022; Ek et al., 2023; Gao & Yin, 2023; Balachandar et al., 2023). Zhang et al. (2021) develop a framework for ranking individualized treatment policies under Rosenbaum’s $\Gamma$ sensitivity model. Namkoong et al. (2020) also use Rosenbaum’s $\Gamma$ sensitivity model to partially identify $V(\pi)$ in the sequential setting. Kallus & Zhou (2021, 2018) develop an approach for learning minimax optimal decision-policies under the MSM. In principle, confounding-robust policy learning frameworks such as these can also be used to recover $\delta$-regret intervals against the confounded data-generating policy (Hatt et al., 2022; Ek et al., 2023; Gao & Yin, 2023). However, these techniques (1) target the mean potential outcome rather than predictive performance measures, and (2) are tied to a fixed causal assumption which may not be applicable in all contexts. In contrast, our framework is interoperable with a range of assumptions (e.g., instrumental variable, MSM, proximal variable) studied in prior OPE literature.
Finally, our work relates to literature studying comparison of human versus algorithmic decision-making policies. Kleinberg et al. (2018) develop a framework for comparing human decisions and algorithmic predictions under selective labels and unobservables. The contraction technique underpinning this approach compares the failure rate of judicial decisions against a predictive model under an assumption that judges have heterogeneous selection rates (Lakkaraju et al., 2017). Later developments formalize this approach into an instrumental variable framework (Chen et al., 2023; Rambachan et al., 2021, 2022). Rambachan et al. (2021) develop a framework for identifying systematic prediction mistakes in historical human decisions. This framework enables utility-based comparisons of human versus algorithmic decision-making policies under varying sets of econometric assumptions. Most recently, Ben-Michael et al. (2024) devise an experimental framework for evaluating human+algorithm hybrid workflows against the quality of decisions that a human or algorithm would make alone. This framework supports post-deployment policy comparisons following randomized assignment to a treatment (algorithmic recommendation) or control (no algorithmic recommendation) condition. In contrast, our framework is designed to support pre-deployment policy comparisons given observational data collected under a status quo policy.
@@ 3 Preliminaries
Let $\pi_{0}:\mathcal{X}\times\mathcal{U}\rightarrow\mathcal{A}$ be a status quo decision-making policy assigning binary actions given measured covariates $X\in\mathcal{X}$ and unmeasured confounders $U\in\mathcal{U}$. Let $\pi:\mathcal{X}\rightarrow\mathcal{A}$ be a proposed policy that assigns binary actions only via covariates. [^4] Our framework is compatible with new policies which are stochastic and deterministic. However we assume that $\pi$ is known. For example, in a medical testing context, $\pi_{0}$ is an existing (e.g., physician based) testing policy, while $\pi$ is a proposed algorithmic policy. Let $D^{\pi_{0}}$, $T^{\pi}$ be random variables indicating actions selected under the status quo and proposed policies, respectively. [^5] We sometimes omit policy superscripts over random variables to ease notation. Let $Y(1)\in\mathcal{Y}$ be the outcome of interest to the policy comparison. This is the potential outcome which would be observed given a positive decision ($D^{\pi_{0}}=1$) selected under the status quo policy (Rubin, 2005). For example, $Y(1)$ denotes the disease status of a patient, or repayment if a lendee is granted a loan. Let $Y\in\mathcal{Y}$ be the binary outcome observed in observational data. In our selective labels setting (Lakkaraju et al., 2017), $Y$ is only observed when an instance received a positive decision ($D^{\pi_{0}}=1$) under $\pi_{0}$. We make the following standard consistency assumption on $Y$ (Rubin, 2005).
@@ 3.1 Problem Formulation
3.2 Partial Identification of Policy Performance
@@ 4 Regret Bound Identification
We now introduce our novel approach for comparing policies under confounding. Our approach yields informative regret intervals by eliminating redundant uncertainty that does not contribute to differential policy performance. We first focus on asymptotic performance bounds, and provide finite-sample analyses in Section 6.
@@
Because $v_{0}(0,0)\in\mathbf{v}_{0}$ cancels when taking a difference across polices, this term does not contribute uncertainty to the $\delta$-regret interval, as it would with the baseline interval. [^8] We provide additional intuition for the benefits of our uncertainty cancellation approach in Appendix E. We now provide a theoretical result characterizing the improvement offered by the $\delta$-interval more generally. We let $I_{\delta}(m,\mathcal{V})=\overline{R}_{\delta}(\pi,\pi_{0};m,\mathcal{V})-\underline{R}_{\delta}(\pi,\pi_{0};m,\mathcal{V})$ and $I(m,\mathcal{V})=\overline{R}(\pi,\pi_{0};m,\mathcal{V})-\underline{R}(\pi,\pi_{0};m,\mathcal{V})$ be the length of the $\delta$-regret and baseline regret intervals, respectively.
@@ 5 Mapping Causal Assumptions to Informative Regret Bounds
The regret interval we recover without imposing additional causal assumptions on the data generating policy tends to be uninformative. Therefore, we now introduce a pointwise bounding functions assumption which we use to tighten regret intervals by shrinking the uncertainty set around partially-identified $v$-statistics. We show that this assumption is implied by a range of causal assumptions studied in prior OPE literature (e.g., MSM, Rosenbaum’s $\Gamma$, IV) in Appendix B. Thus, our $\delta$-regret interval yields tighter uncertainty quantification than the baseline interval for any causal assumption implying pointwise bounding functions (Figure 2).
@@
We only use intervals $\mathcal{H}(v_{1}(t,0);\tau)$ for this set definition because an analogous bound is implied over $v_{0}(t,0)$ by the constraint requiring that $\sum_{y}v_{y}(t,0)=\rho_{t0}$. We quantify the size of $\mathcal{V}(p;\tau)$ via the Lebesgue measure, which is defined as the cartesion product over partially identified intervals $\lambda(\mathcal{V}(p;\tau))=|\overline{v}_{1}(1,0)-\underline{v}_{1}(1,0)||\overline{v}_{1}(0,0)-\underline{v}_{1}(0,0)|$. We now show that $\mathcal{V}(p;\tau)$ is the smallest uncertainty set one could construct based on the observational distribution and a pair of bounding functions.
@@
Algorithm 1 Plug-in regret bound estimator
  for $k=1$ to $K$ do
     Learn $\hat{\eta}_{k}=(\hat{e}_{1,k},\hat{\tau}_{k})$ using $\mathcal{O}_{-k}$
     Construct $\hat{\mathcal{V}}_{k}(p;\hat{\tau})$
  end for